I have two question that I don't understand how to do. Please help
1. find the exact value of: tan(-3pi)
2. Find the exact value of: sin 60 degrees - cos60 degrees.
I have to all my work for both problems. Will someone please help.
Thanks

Answer by stanbon(75887)

1. find the exact value of: tan(-3pi) = 0
2. Find the exact value of: sin 60 degrees - cos60 degrees.
= (sqrt(3)/2) - (1/2) = (sqrt(3)-1)/2
========================
Cheers,
Stan H.